### 22 Mar 2024 v2

# Gen-T: Table Reclamation in Data Lakes

Abstract
We introduce the problem of Table Reclamation. Given a Source Table and a large table repository, reclamation finds a set of tables that, when integrated, reproduce the source table as closely as possible. Unlike query discovery problems like Query-by-Example or by-Target, Table Reclamation focuses on reclaiming the data in the Source Table as fully as possible using real tables that may be incomplete or inconsistent. To do this, we define a new measure of table similarity, called error-aware instance similarity, to measure how close a reclaimed table is to a Source Table, a measure grounded in instance similarity used in data exchange. Our search covers not only Select-Project-Join queries, but integration queries with unions, outerjoins, and the unary operators subsumption and complementation that have been shown to be important in data integration and fusion. Using reclamation, a data scientist can understand if any tables in a repository can be used to exactly reclaim a tuple in the Source. If not, one can understand if this is due to differences in values or to incompleteness in the data. Our solution, Gen-T, performs table discovery to retrieve a set of candidate tables from the table repository, filters these down to a set of originating tables, then integrates these tables to reclaim the Source as closely as possible. We show that our solution, while approximate, is accurate, efficient and scalable in the size of the table repository with experiments on real data lakes containing up to 15K tables, where the average number of tuples varies from small (web tables) to extremely large (open data tables) up to 1M tuples.
I Introduction
We introduce the problem of Table Reclamation where we are given a source table and seek to find a set of tables from a data lake (a large table repository) that, when integrated, reproduce the source table as closely as possible. We begin with an example showing how table reclamation can be used.
[…]
Suppose a user is reading a news article that reports the demographics of employees in Top US tech companies in 2021 (top blue table in Figure 1). The user has access to 2021 Microsoft’s Diversity Report [1] which seems to contradict the numbers in the news article (bottom green table in Figure 1). Using table reclamation and her data lake (which may include wikitables, NYTimes Data and other public and private datasets), the user can ask if there is a set of tables that, when integrated, recreates the data in the news article (blue table). Table reclamation is able to reproduce this table using a number of tables including World_MS_Ethnicity and World_MS_Employees (which, after being joined and a selection on 2021 is applied, produce the first tuple). We call the tables used in reclamation, originating tables. Other tables can be unioned with the Microsoft tables to reclaim the other tuples. From this (the originating tables including their meta-data and data), a user can understand that while her table (green) is reporting US statistics numbers, the article is reporting international numbers.
Unlike the well-known problem of Data Provenance [2, 3], we do not have prior knowledge of the query or tables that were originally used to create a Source Table. Instead, we focus on recovering possible tables that, when integrated, confirm the data values and facts in a Source Table. Table reclamation is related to the common Query-By-Example (QBE) or Query-By-Target (QBT) that discover a query over input tables that produces an instance-equivalent table to the given example output table [4, 5, 6, 7, 8, 9]. In order to generalize to a data lake setting, we do not assume we know a complete and correct set of input tables. Rather, we use an additional step of finding candidate tables within or across data lakes that may contribute to the Source Table (i.e., that may be originating tables). Also, existing QBE/QBT systems focus primarily on discovering (Select)-Project-Join queries over (largely complete) relational tables [10, 11, 12, 13, 14], with some using both the data values and the schema of the tables. Due to the noise and heterogeneity of data lake tables, these queries may not be sufficient to fully integrate data lake tables to produce a given Source table. So, we aim to recover Select-Project-Join-Union queries using only the data values, since the metadata of data lake tables may be missing or inconsistent [15, 16, 17, 18]. We also consider operations that have proven to be important in data integration and data fusion of incomplete data, namely subsumption and complementation [19].
Our goal is to reclaim the source table completely, but this may not always be possible.
[…]
Continuing our example, it is possible that the best reclamation we can find has null values for the percentages of Hispanic employees for Google and a different number of Asian employees (20% instead of 24%). These differences indicate that the source data about Google was not completely found within the data lake. The user can analyze the originating tables returned by our approach to understand these differences. As an example, it may be that the originating tables for the Google data in the repository are European in origin and do not report values for all categories like Hispanic employees as this is a protected category under US, but not European law.
If certain tuples cannot be reclaimed, a data scientist would know these are not derivable from her data lake. If the reclaimed table contains different values from the source, a scientist can investigate whether the source values are wrong or if they are valid corrections to errors in the originating tables. Unlike traditional QBE and QBT approaches, our focus is on the data (rather than the query) and on understanding what data in a Source Table (and only data in a Source Table) can be reclaimed. We make the following contributions.
We define the novel problem of Table Reclamation – finding a set of originating tables that, when integrated, can reproduce a Source Table as closely as possible.
To evaluate how close a reclaimed table is to a Source Table, we define a new error-aware instance similarity (EIS) score that is a principled extension of instance similarity used in data exchange [20]), and show how it can be computed efficiently.
We present an approximate Table Reclamation solution named Gen-T that performs table discovery to retrieve a set of candidate tables, and filters out poor candidates using a novel table representation that simulates table integration without performing expensive integration operations. The remaining originating tables are integrated to produce a table whose values are as close as possible to the Source Table.
We conduct extensive experiments on real and synthetic data lakes, showing that Gen-T outperforms all baseline methods. Gen-T reclaims 5X more values from Source Tables than the best-performing baseline. We perform an ablation study on the sensitivity of Gen-T to erroneous data (data that cannot be reclaimed) and to incomplete data.
We show that our solution is efficient and scalable to the size of the data lake with experiments on real data lakes containing up to 15K tables, where the average table size (number of tuples) varies from small (web tables) to extremely large (open data tables) with on average over 1M tuples. In addition, our solution is scalable to large source tables, with experiments on source tables containing up to 22 columns and 1K rows.
II Overview
A data scientist provides a Source Table that she would like to reclaim by understanding if it can be produced by integrating any combination of tables within a data lake. Specifically, we aim to determine a set of tables from which the Source Table’s values may originate (termed originating tables), and use them to reclaim (regenerate) the Source Table. Given our data lake setting where tables can be changed autonomously, we formulate the problem as an approximate search of finding a set of tables that can best be used to reclaim the Source, as closely as possible.
Unlike many existing Query-by-Example [5, 10, 11, 12, 13, 14, 6, 7, 8] or by-target [9] approaches, we do not assume that we know the exact set of input tables whose values first formed the Source Table or even if the Source Table can be reclaimed. In addition, while we do not assume tables in a data lake to have keys or any foreign key relationships, we assume the Source Table to have a (possibly multi-attribute) key, which can be found using existing mining techniques [21, 22]. This is a restriction, but it is made to make the instance comparison (which is done often in the algorithm) efficient. Without the source table having a key, instance similarity requires homomorphism checks which is NP-hard [20]. In general, we do not assume that metadata is available for any tables (column names are included in examples only for clarity).
To solve the problem of table reclamation, we use a two-step solution. First, we discover tables from the data lake that share values with the Source Table and therefore may have created portions of it , we call these candidate tables. Then, we search for ways of combining subsets of these tables to regenerate the Source Table.
Figure 2 shows the pipeline of Gen-T. The input is a Source Table and the output a Reclaimed Source Table and its originating tables. In the Table Discovery phase (Section V-A), Gen-T discovers a set of candidate tables whose values may have contributed to the creation of the Source Table. Then, we apply our novel solution of representing tables as matrices in order to simulate table integration via matrix traversal (Section V-A2, V-A3). The goal of this step is to refine the set of candidate tables to a set of originating tables, and essentially filter out candidate tables that are not needed before performing table integration. To efficiently retrieve a set of candidate tables, Gen-T uses an existing, data-driven table discovery method that has no guarantees for this problem setting. Gen-T then prunes the candidate tables to a set of originating tables by computing each candidate table’s similarity with a given Source Table and simulating table integration. Once Matrix Traversal pinpoints a set of originating tables, we integrate these originating tables in the Table Reclamation phase (Section V-B) and produce a reclaimed Source Table.
[…]
Suppose a user has the top left, green table in Figure 3 as a Source Table. To reclaim this table, we use table discovery to find a subset of tables in the data lake with overlapping values – in this example, tables A, B, C, and D. However, Table C contains contradicting non-null values in the “Gender” column compared to values in the same tuples of the “Gender” column in the Source Table. We demonstrate the consequence of directly integrating all these tables, including Table C. The top right, yellow table is the integration result from the state-of-the-art full disjunction (FD) method [23, 24] and the bottom right table shows the result using one possible outerjoin order that may be learned by Auto-Pipeline (a by-target approach) [9]. The resulting tables contain different values in the Gender column (in red) with respect to the corresponding value in the Source Table. These values originate from Table C. When possible, we need to refine the set of candidate tables to filter out tables like Table C that produce integrated tables with erroneous values that do not make the Source Table. Note in this example, the integration of Tables A, B and D alone produces a better reclamation than using all four candidates.
We present related work next in Section III. We define the problem of Table Reclamation and preliminaries for our solution in Section IV. Then, we discuss the two steps in Gen-T– Table Discovery phase (Section V-A) and the Table Reclamation phase (Section V-B). Finally, the experiments in Section VI show the effectiveness, scalability, and generalizability of Gen-T and we conclude with open problems and exciting directions for this new area (Section VII).
III Related Work
We now discuss related work on Table Discovery and Integration and work related to finding the origins of tables.
Table Discovery: Table Discovery has a rich literature, specifically keyword search over tables, unionable table search, and joinable table search. Early work such as Octopus [25] and Google Dataset Search [26], support keyword search over the metadata of tables [16, 27] and smaller scale web-tables [28, 29]. Data-driven table discovery systems [30, 31, 32, 33, 15] were then developed to find schema complements, entity complements, joinable tables, and unionable tables.
For joinable table search, early systems use schema matching or syntactic similarities between tables’ metadata, such as Jaccard similarity [34, 35]. LSH Ensemble [31] makes use of approximate set containment between column values and supports set-containment search using LSH indexing. JOSIE [32] uses exact set containment to retrieve joinable tables that can be equi-joined with a column in the user’s table. MATE [36] supports multi-attribute join with a user’s table. DeepJoin [37] leverages a deep learning model to retrieve equi-joinable and semantically joinable tables. These systems can be used to retrieve a set of candidate tables that have high set similarity with a given user’s table.
For table union search, early systems also used schema similarity [30, 38]. Using data (rather than metadata), a formal problem statement for unionability was first defined by Nargesian et al. [15] who presented a data-driven solution that leverages syntactic, semantic, and natural language measures. This problem was refined by SANTOS [39] to consider relationship semantics as well as column semantics. Most recently, Starmie [40] offers a scalable solution to finding unionable tables that leverages the entire table context to encode its semantics. Although our method also retrieves relevant tables to a user’s table, we aim to retrieve tables for a specific task – reclaiming the user’s table. Finally, other recent work [41] presents a goal-oriented discovery for specific downstream tasks, aiming to augment columns. We tailor table discovery towards the goal of reclaiming the Source Table.
Table Integration: Lehmberg et al. [42] stitches unionable tables together, but does not support join augmentation of tables. Recently, ALITE [24] performs full disjunction (FD) [23] to maximally combine tuples from a set of tables (intuitively, full disjunction is a commutative and associative form of full outer join). Our goal is to reproduce the given Source Table, which may contain incomplete tuples, so we do not aim to maximally combine tuples if it produces a table that is not identical to the Source Table. Nonetheless, ALITE is a candidate baseline for Gen-T, as it offers a state-of-the-art integration solution.
Preceding the table integration process, there are pre-integration tasks to find alignments between table elements. First, instance-based schema matching determines how the schemas of two tables align to prepare for integration [43, 44, 45, 46, 47, 48]. Our solution, Gen-T, aligns schemas implicitly by renaming columns in the retrieved tables with the column of the source table that best matches. Entity matching [49, 50, 51, 52, 53, 54, 55, 56, 57] is another common pre-integration task, aiming to align tuples for cleaning or joining tables. In our context, since we assume that the Source Table has a key, tuples can be aligned by matching using equality on the key.
Finding Origins of Tables: Our problem setting of tracing a Source Table’s values back to its origins can be related to Data Provenance [2, 3], which given a query and its output table, explains from where the (values or) tuples originate, why and how they were produced. However, in our problem setting, we do not know the query or originating tables that were originally used to create the Source Table.
Query-By-Example (QBE) is a popular approach. The original QBE was a language allowing nonexpert users to query a database [58]. More recently this term has been used for methods that are given a pair of matching input and output tables, and the task is to synthesize a query from the input to the output[5]. For this task, some systems only consider Project and Join operators [10, 11, 12, 59], whereas others also consider the Select operator [13, 14]. Others output a set of queries that could reproduce the example output table, given the input table [6, 7]. AutoPandas [8] performs transformation-by-example by synthesizing Pandas programs rather than SQL queries. More recently, proposed techniques relax the assumption that a set of tables from which the query table can be generated is provided [59, 60, 61]. Instead, they discover a set of tables that, when integrated, produce a table that contains the query table. These methods often expect only a partial query table with a small set of attributes and possibly a set of tuple examples (for example, Ver [59] uses queries that are tables of 2 columns and 3 rows). Their goal is to generate an output table that completes this query table by returning a table that contains many additional tuples in addition to those in the query table. However in our problem, our goal is to reproduce all and only tuples from the source table. Nonetheless, since Ver [59] is the state-of-the-art Query-by-Example method, we use it as a baseline for Gen-T.
Auto-Pipeline [9] defines Query-By-Target, with the similar goal of synthesizing the pipeline used to create the target table, given the target table and a set of input tables. Using the synthesized pipeline on the input tables, it then produces a table that “schematically” aligns with the input target table. As the state-of-the-art in this line of work, it is a baseline for our approach. In both By-Example and By-Target paradigms, many systems assume that the set of input tables on which the system synthesizes a query to generate the example or target table is known and perhaps more importantly are known to contain the tuples and columns needed to reproduce the output table. In our problem, we do not assume this is the case.
IV Problem Definition and Preliminaries
We first discuss how we evaluate a possible reclaimed Source Table and then define the problem of Table Reclamation. Then, we describe preliminaries for our solution, specifically the set of table operators with which we perform table integration to produce a possible reclaimed Source Table.
IV-A EIS Score and Problem Definition
To evaluate a possible reclaimed table, we compare it with the Source Table to see how close they are. The problem of comparing database instances is prevalent in many applications such as analyzing how a dataset has evolved over time (e.g., data versioning) [62], evaluating data cleaning solutions (e.g., compare a clean instance produced by a data repair algorithm against a gold standard) [63], or comparing solutions generated by data exchange or transformation systems [64, 20]. This similarity score requires the computation of a mapping between the tuples across instances, which can be used to explain the result. The most general measures rely on homomorphism checking and are NP-hard [65]. Since data lake tables do not have keys, integrating them can produce multiple copies of tuples from the source table. We will align data lake tuples with a single source tuple where the lake and source tuple share the same key value (i.e., are aligned tuples iff they share the same values on key attributes). Hence, multiple lake tuples may align with the same source tuple, and some will align with no source tuple. But a lake tuple will align with at most one source tuple and because of this, we can do the mapping efficiently.
We propose an error-aware instance similarity score as a generalization of instance similarity defined by Alexe et al. [20], which has been widely used in the integration literature [66, 67, 68, 69]. Instance similarity quantifies the preservation of data associations when source data is exchanged into target database [20]. It relies on computing homomorphisms (as it is designed for data exchange where keys are not assumed). Considering two relations with the same schema, if two tuples are mapped, they define the tuple similarity as the ratio of the number of values that are shared over the size (cardinality) of the tuples. Note that in data exchange two tuples cannot be mapped if they disagree on any non-null values. In our setting, we can map tuples that differ on their non-null values. Hence, we define an error-aware tuple similarity that penalizes mismatching (erroneous) values.
[…]
Given two tuples and with the same schema containing non-key attributes, where and share the same key value. Let be the number of non-key attributes on which and share the same value and be the number of non-key attributes on which and have different values and is a non-null value. Then the error-aware tuple similarity is111Note the tuple similarity defined by Alexe et al. [20] is :
Since a tuple can map to more than one tuple, Alexe et al. define instance similarity using the maximum tuple similarity score and we do the same in defining error-aware instance similarity, but we use error-aware tuple similarity rather than tuple similarity. For completeness (and since in our experiments we use both measures), we now define both.
[…]
Let be a source table (with key attributes ) and a possible reclaimed table with the same schema that has non-key attributes. Note that does not have to satisfy the key constraint. For a tuple , let . Then the instance similarity of and is:
The Error-Aware Instance Similarity (EIS) of and is (for a normalized score in range[0,1]):
We illustrate these scores in an example based on Figure 3.
[…]
Consider Source Table (with key column “ID”) from Figure 3 (top-left green table), and two possible reclaimed tables, (top-right yellow table) and (bottom-right yellow table). Their aligned tuples are shown in Figure 4. Notice the instance similarity score (counting only matching values) of and is higher than of and .
However, we want to favor that contains nullified (unknown) values (one tuple matching the source correctly and two tuples missing values), over , which has reclaimed a possibly erroneous value for a source null. Using EIS score, has a higher similarity with than .
Using EIS Score as a similarity measure to compare a possible reclaimed table and Source Table , we aim to solve the following problem:
[…]
Given a collection of tables and a Source Table , find a set of originating tables such that its integration produces a reclaimed table with the maximum EIS Score to .
IV-B Preliminaries: Integration Operators
We now present the operators we use to produce a possible reclaimed Source Table. We will show that they are sufficient for integration, inspired by recent work on data lake integration [24]. First the unary operators.
Projection(): Project on specified columns of the table.
Selection(): Select tuples that satisfy a specified condition.
Subsumption() [23]: Given tuples
with the same schema, subsumes if for every attribute on which they are both non-null they have the same value, and
contains one or more attributes with a non-null value where has nulls. Applying subsumption we remove . Applying on a table involves repeatedly applying subsumption and discarding the subsumed tuples ().
Complementation() [19, 70]: Given tuples
with the same schema, complements if they share at least one non-null column value, and contains some non-null values where has nulls while contains some non-null values where has nulls. The tuples must agree on all values on which they are both non-null.
Applying on and produces a single tuple that contains all non-null values of either (both) tuples and is null only if both and are null. Applying on a table produces a table with no complementing tuples and involves repeatedly applying complementation to pairs of tuples.
We use a single binary operator natural Outer Union, and we assume the schemas of the tables have been aligned so unionable columns share the same name [71].
Outer Union() [72]: Union two tables, even if their schemas are not equal. The result contains the union of the columns from both tables. If a column is missing from one table (), but appears in the other table (), then in the result, the tuples of contain a null () in their column. This operator is commutative and associative. Note that when applied to tables with the same schema, outer union is the same as inner union.
To make reclamation search more efficient, we use the fact that Outer Union and the set of unary operators above can be used to represent any SPJU query. Using this result, our search will focus on outer union and the unary operators.
Theorem 1 (Representative Operators)
Given two tables that contain no duplicate tuples, and no tuples that can be subsumed or complemented, for all SPJU queries, there exists an equivalent query consisting of only Outer Union and the four unary operators (selection, projection, complementation, and subsumption).222The proof is included in our technical report [73].
V Table Reclamation using Gen-T
We describe Table Discovery (Section V-A), which finds a set of originating tables that we can effectively integrate to reclaim a source table (Section V-B).
V-A Table Discovery
We first discover a set of candidate tables, after which we discuss a novel methodology, termed Matrix Traversal, to refine this set into a set of originating tables.
V-A1 Candidate Table Retrieval
Discovering a set of candidate tables requires discovering tables that share some of the same values as the Source Table in an efficient manner. In the context of data lakes, where metadata is inconsistent or missing, searching using schema names is unreliable [16, 17, 18, 74]. However, we can use any existing data-driven table discovery approach that is scalable in a data lake setting.
With a set of top- tables returned as relevant to the Source Table, we need to verify the set similarities of their values with the Source Table. To do so, we retrieve candidate tables among the previously discovered tables using a set similarity algorithm. This could be done efficiently with a system like JOSIE [32] that computes exact set containment or MATE [36] that supports multi-attribute joins. In addition to finding candidate tables containing columns that have high set similarity with a Source Table, we also diversify the set of candidate tables such that each candidate table has minimal overlap with other candidates. This is especially important in public data lakes, which tend to have multiple versions of the same tables [62, 75] and a large percentage of duplicate column sets [32]. By diversifying candidates, each candidate may overlap with different values in , as illustrated in the following example.
[…]
Suppose we have the Source Table from Figure 3. In addition to data lake tables A, B, C, D, we also have Table E, an exact duplicate of Table D. If we only rank these tables using set overlap with the Source Table, Tables D and (its duplicate) Table E become top candidates, since all their columns have high set overlap with those in the Source Table. However, Table E does not add any new information when integrated with Table D. Thus, diversifying the set of candidate tables decreases Table E’s score, pushing other tables such as Table A higher in the top- ranking.
With a diverse set of candidates found for each column in a Source Table , we ensure that each candidate table still has high set overlap with the Source Table across related columns. To do so, we find all tuples in a candidate table that share column values with . We verify that for each column that has high set overlap with a column in , it still has high set overlap within these tuples. For columns from the candidate table that have high value overlap with columns in , we rename them with names of corresponding columns in . This way, we implicitly perform schema matching between each candidate table and . Finally, we check for and remove any candidate table whose columns and column values are subsumed by other candidate tables.
V-A2 Matrix Traversal
With a set of diversified candidate tables, we could potentially enter the table integration phase (Section V-B). However, it may be computationally expensive to directly integrate all candidate tables. Thus, we need to refine the set of candidate tables to a set of originating tables containing a maximum set of aligned tuples with respect to the Source Table. To do so, we emulate the table integration process and see what candidate tables are necessary to reclaim our Source Table. By simulating tuple alignment, we can uncover erroneous aligned tuples with respect to the Source Table, and discard tables that could decrease the EIS Score.
First, we need to align tuples in candidate tables to tuples in a Source Table, based on shared values with the key attribute from the Source Table. To do so, we need to ensure that each candidate table contains a key column of the Source Table. If it does not, we greedily find a best way to join it with candidates that include the source key. We use standard join cardinality estimation to find a path that covers the most source key values and is as close to functional as possible (this procedure is denoted as Expand()).333The Expand algorithm is included in the technical report [73]. This way, all tables can align its tuples with the Source Table using key values.
We represent aligned tuples and shared columns from a candidate table in the form of a matrix. To encode aligned tuples, we initialize the matrices to have the same dimensions as the Source Table , such that the matrix indices represent the Source Table’s indices. For each key value and its associated column values in the Source Table, check if the value appears in the candidate table at the corresponding column and key value. If so, then the matrix has 1 at the same index as the value’s index in Source Table, and 0 otherwise.
Next, we simulate table integration by applying the logical OR on the matrices, which takes the maximum value at each position. Suppose we have tuples from matrices , respectively, at the same row index . We want to combine values from at column . Assuming that , the produced tuple contains the following value at position (): . If is 1 and is 0, for example, is 1, which is equivalent to the logical OR. This is comparable to applying the Outer Union () of two tables, and Subsumption () and Complementation () on the resulting table (refer to Theorem 1). In this table integration, for tuples that share a non-null value at the same column, the resulting tuple is formed such that for every column , . Thus, both table and matrix integrations maximally combine tuples such that non-null values replace a null value at the same index.
We demonstrate matrix initialization and traversal in Algorithm 1. Given a set of candidate tables and Source Table , we first ensure that each candidate table contains a key column from by joining candidate tables that do not share a key column with with those that do (Expand() on line 1). This way, we align tuples in each candidate table with respect to and initialize each candidate table’s matrix representation (Line 1). Then, we traverse over the matrices and perform the logical OR operator to combine a pair of matrices in (Line 1). To evaluate the resulting matrix, we check the fraction of 1’s in the matrix, which represent the number of values in the resulting table integration found in the Source Table, thus evaluating the EIS score. At each step of the matrix traversal (Line 1), including the start (Lines 1-1), we choose the matrix that results in a matrix containing the most 1’s in . This traversal ends when either all matrices have been traversed (Line 1), or the percentage of 1’s in the resulting matrix converges (Lines 1-1). We return the set of tables used in the final traversal as the set of originating tables to perform table integration. If a candidate table that was joined with other candidates to contain a key column from (from Expand() on line 1) becomes an originating table, we include its expanded form in the returned set.
V-A3 Three-Valued Matrices
Previously, we use matrices populated with binary values to represent aligned tuples with respect to the Source Table. However, this representation cannot distinguish between nullified and erroneous aligned tuples with respect to the Source Table. Specifically, it does not account for cases in which a tuple in the Source Table and an aligned tuple in a candidate table have different non-null values in the same column, and if a tuple in the Source Table has a null value while the aligned tuple has a non-null value at the same column. Rather, it represents both types of values as 0 in the matrices. In actuality, when we apply Outer Union on two tables with aligned tuples containing different non-null values in the same column, we keep the tuples separate. Thus, we need to distinguish between nullified and erroneous aligned tuples in the matrix representation (Line 1 in Algorithm 1). To do so, we use three-valued matrices, in which we encode a 1 if a candidate table shares the same value with the Source Table at the same index in an aligned tuple, 0 if a candidate table contains a null where the Source Table has a non-null value at the same index, and -1 if a candidate table contains a non-null value that contradicts with the Source Table’s value at the same index (shown in Figure 5). Formally, given Source Table and candidate table , we populate position for each aligned tuple in matrix as:
Combining the amended three-valued matrix representations with the logical OR takes the maximum of two truth-values at each index. Specifically, if we have two tuples from two matrices that contain a 1 and -1 at the same position, applying logical OR would choose the 1 [76]. However, in practice when applying Outer Union on two tuples with contradicting non-null values at the same index, the resulting integration would contain both tuples. Thus, we need to keep both tuples from the matrices if they contain different non-0 values at the same index. We re-define () (Line 1) between two matrices, given tuples at the same row index accordingly.
This way, we keep tuples separated if they contain contradicting, non-0 values at the same position. Else we apply logical OR and take the maximum of truth values element-wise. The new Combine() function is illustrated in Example V-A3.
The new () could result in matrices with more rows than in the Source Table. Thus, we encode each matrix as a dictionary, with each key value in the Source Table as a dictionary key, and the list of aligned tuples in the resulting matrix with respect to a tuple in the Source Table as values.
[…]
Given the Source Table from Figure 3, Figure 5 shows the result of integrating tables A, B, and C and their matrix representations. We start with matrix A with the largest number of correct values. Integrating matrices A and B produces more correct values after taking logical OR’s of 0’s and 1’s (from the Combine() function). When combining its resulting matrix with matrix C, we find a (1) and () in the first tuple for column “Gender”. In this case, we keep both tuples from OR(A, B) and C. For all other value-pairs, we take the logical OR. When we integrate tables A, B, C, we also find tuples for “Name”=Smith to be separate, since we check if we over-combine tuples and replace correct nulls (Algorithm 2). As a result, the matrix representation of this table integration result is exactly the output of the matrix integration, and so table integration is simulated by integration of matrix representations.
For of the start (Lines 1-1) and resulting matrices (Line 1), we evaluate the EIS Score by taking the aligned tuple with the largest number of aligned values to its corresponding tuple in the Source Table, or the largest number of 1’s. To find the EIS score (Equation 3) between a tuple in a resulting matrix that shares a key value with a Source tuple , we set to be the number of non-key attributes for which tuple has (1)’s, representing shared values between and . For , we take the number of non-key attributes for which has (-1)’s, representing different, non-null values. Thus, we treat correct, nullified, and erroneous aligned tuples with respect to the Source Table in different manners, and combine their matrix representations depending on the behavior of applying Outer Union and unary operators.
V-B Table Reclamation via Integration
With the set of discovered originating tables, we now integrate them to reproduce the Source Table as closely as possible. Using the set of representative operators (Theorem 1), we present an efficient algorithm to explore the integration space to reclaim the Source Table.
Given a set of originating tables (), a Source Table () and using , Gen-T outputs a table () that reclaims as best as possible. Our table integration method is depicted in Algorithm 2. From Section V-A1, we ensure that all candidate tables (and thus originating tables) contain ’s columns.
V-B1 Preprocessing
First, we project out columns not in (), and select tuples whose values are in the source key column (). Hence, we only keep columns and tuples that overlap with (ProjectSelect(), Line 2). We then union all originating tables that share the same schema (InnerUnion(), Line 2) to reduce the space of tables we need to explore. To prevent over-combining tuples that share nulls with tuples in , for each table , we find tuples in that share key values and contain nulls in the same columns, and replace these nulls in with unique labeled non-null values (LabelSourceNulls(), Line 2). Finally, we remove duplicate tuples, subsumed tuples (), and take the resulting tuples of complementation () (TakeMinimalForm() on Line 2).
V-B2 Integration
We integrate all resulting tables , all of which contain the source key column. At each iteration of , we outer union with the integrated result so far, . Next, we check if applying Complementation() and Subsumption() on results in a table that has a higher EIS score in the evaluateSimilarity() function. This lets us check if these operators are over-combining tuples (e.g., removing a subsumed tuple that is identical to a tuple in ) and decreasing the number of values shared with . After iterating through all tables from the input set, we revert the previous labeling of shared nulls with (RemoveLabeledNulls() on Line 2). To ensure that the resulting integrated table () has the same schema as , we add null columns in for every column it does not share with (Line 2). Finally, we return the resulting integration as a possible reclaimed table.
VI Experiments
We evaluate Gen-T on benchmarks with tables containing real instances from well-known Benchmarks [77, 78] along with tables from real data lakes [39, 79]. Our baselines are modified related techniques to solve the reclamation problem. Section VI-B shows that Gen-T can perfectly reclaim 15-17 Source Tables, whereas most baselines only fully reclaim 1-3 Source Table across benchmarks. Section VI-C shows that Gen-T is 5X faster than the next-fastest baseline on a large data lake. Finally, Section VI-D shows that Gen-T generalizes to a different real-world application.
VI-A Experimental Setup
We implement Gen-T in Python on a CentOS server with Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz processor. Code for Gen-T and all the baselines is publicly available [80]. We evaluate Gen-T on 6 benchmarks (see Table I).
| Benchmark | # Tables | # Cols | Avg Rows | Size (MB) |
| TP-TR Small | 32 | 244 | 782 | 3 |
| TP-TR Med | 32 | 244 | 10.8K | 40 |
| TP-TR Large | 32 | 244 | 1M | 3.9K |
| SANTOS Large +TP-TR Med | 11K | 122K | 7.7K | 11K |
| T2D Gold | 515 | 2,147 | 74 | 4 |
| WDC Sample +T2D Gold | 15K | 75K | 14 | 66 |
TP-TR Benchmarks: We use the 8 tables from the TPC-H benchmark [77], which contain business information including customers, products, suppliers, nations, etc. Using these tables, we create three versions of a benchmark suite titled TP-Table Reclamation (TP-TR). TP-TR Large has TPC-H tables with original table sizes. TP-TR Med has TPC-H tables that are each 1/100 of its original table’s rows, and TP-TR Small has TPC-H tables that are each 1/1000 of its original table’s rows. In each, we take each of the 8 tables and create 4 versions of the same table – creating 32 tables in total. For two versions, we randomly nullify different subsets of values, and for the other two versions, we randomly inject different non-null (erroneous) values in different subsets of values. For the majority of the experiments, the number of nulls (respectively, erroneous values) is 50% meaning that we randomly take 50% of each table’s values and replace them with nulls (respectively, replace them with different new strings). In an ablation study (Section VI-B last two paragraphs), we vary (independently) the number of nulls/erroneous values from 10% to 90%. Our goal in the table discovery phase is then to filter out originating tables with injected non-null noise, so that resulting reclaimed tables do not contain erroneous values. Thus, we seek to verify that our approach uses the nullified versions rather than the erroneous versions so that combining them can reproduce the Source Table.
The Source Tables for the TP-TR benchmarks are created using all 8 of the original (unmodified) TPC-H tables over which we randomly generated 26 queries each having a subset of operators {}. In these 26 queries, the number of operations ranges from 2 (just ), to 9, such that the query with the maximum number of unions contains 4 unioned tables, and the query with the maximum number of joins joins 3 tables. We ran the same queries on each TP-TR benchmark to create 26 Source Tables for the TP-TR Small benchmark containing an average of 9 columns and 27 rows, and 26 Source Tables for the TP-TR Med and TP-TR Large benchmarks that have an average of 9 columns and 1K rows.
SANTOS Large +TP-TR Med Benchmark: In addition, to further assess the effectiveness and scalability of our table discovery method, we embed TP-TR Med into a real, large data lake SANTOS Large [39]. In doing so, we evaluate how well Gen-T prunes a potentially large set of candidate tables retrieved from a large data lake to a smaller set of originating tables that can more accurately reclaim a Source Table when integrated. We use the same Source Tables as for TP-TR Med.
T2D Gold Benchmark: In addition, we explore the real-world application of our method with the T2D Gold Benchmark [78], which takes web tables and matches them to properties from DBpedia. This benchmark was not originally created for the problem of Table Reclamation, so we test the generalizability of Gen-T by seeing if it can reclaim any of this benchmark’s tables. We take 515 raw tables that contain some non-numerical columns and a key column. We do not have prior knowledge of whether or not any of these 515 tables can be “reclaimed” as a Source. Thus, we iterate through each of the 515 tables as potential sources.
WDC Sample +T2D Gold: To further assess the effectiveness of Gen-T, we embed T2D Gold tables into a sample of the WDC web table corpus [79], which contains 15K relational web tables. This way, we can examine how well Gen-T prunes a large set of candidate tables found from a large table corpus to a small set of originating tables that can be integrated to reclaim a Source Table.
VI-A1 Baselines
We compare Gen-T to the state-of-the-art for by-target synthesis, Auto-Pipeline [9], the state-of-the-art for Query-by-example, Ver [59], and the state-of-the-art for table integration, ALITE [24] which were modified for the reclamation problem.
Auto-Pipeline has a similar framework to our problem in discovering the integration (query or pipeline) that reclaims a Source Table. However, Gen-T does not assume to have the perfect set of input tables from which we can synthesize the query that reproduces the Source Table. Auto-Pipeline has both query-search and deep reinforcement learning approaches, but since we propose an unsupervised approach, we use the query-search variation as our baseline. Auto-Pipeline’s code implementation is not openly available, so we adopted an open re-implementation of their search approach [62], which adapts the framework in Foofah [81], and, for a fair comparison, revised their set of table operators to only contain table operators that Gen-T considers ({}). We call this re-implemented, adapted baseline Auto-Pipeline*. Since Auto-Pipeline’s benchmarks contain small tables, and most of their operators are string-transformation operators, we do not consider their benchmarks for our experiments.
Ver has a similar objective of discovering and integrating tables to produce a table that contains the source table and other similar tuples. In contrast, Gen-T aims to reproduce only the tuples from the source table. Ver takes small source tables as input (e.g., source tables with 2 columns and 3 rows [59]), so we query Ver with two columns from the Source Tables. We evaluate the output table for each run, and aggregate the results to evaluate the entire source table.
We validate the need for our Matrix Traversal rather than directly integrating the set of candidate tables returned from Set Similarity (Section V-A1), by comparing against ALITE provided with the set of candidate tables from Set Similarity as input. We also compare with a variation of ALITE, which we call ALITE-PS, that, similar to Gen-T, first performs projection and selection to match the Source Table before performing table integration. ALITE without project and select is much slower as it creates a larger integration result.
For all baselines on the TP-TR benchmarks, we create another variant in which we give each method a specific integrating set (int. set) of tables as input, rather than the full set of candidate tables returned from Set Similarity. We know what subset of tables from the 8 original tables were used to create the 26 Source Tables, so, we know that a perfect reclamation contains variants of these tables. Thus, for all original tables used to create each Source Table, the integrating set includes all variations (2 nullified and 2 erroneous versions of each original table) of these tables.
VI-A2 Metrics
For effectiveness, we evaluate how many values in a Source Table have been reclaimed, or how similar the values in the reclaimed table are to those of the Source Table. Thus, the Source Tables are essentially our ground truth in that we see how many of its values we can reproduce. In an aligned tuple in a reclaimed table with respect to a Source Table, it contains an erroneous value if there is a non-null, different value at a given column compared to a value in the same column in the Source Table. Similarly, it contains a nullified value if it contains a null value in a column where the Source tuple contains a non-null value (see Section IV-A).
| TP-TR Med | SANTOS Large +TP-TR Med | TP-TR Large | ||||||||||
| Method | Rec | Pre | Inst-Div. | D | Rec | Pre | Inst-Div. | D | Rec | Pre | Inst-Div. | D |
| ALITE | 0.662 | 0.202 | 0.100 | 35.831 | — | — | — | — | — | — | — | — |
| ALITE w/ int. set | 0.694 | 0.202 | 0.085 | 36.348 | 0.694 | 0.202 | 0.085 | 36.348 | — | — | — | — |
| ALITE-PS | 0.880 | 0.556 | 0.009 | 3.524 | 0.842 | 0.554 | 0.011 | 4.629 | 0.775 | 0.521 | 0.049 | 21.978 |
| ALITE-PS w/ int. set | 0.880 | 0.569 | 0.009 | 3.524 | 0.880 | 0.569 | 0.009 | 3.524 | 0.880 | 0.569 | 0.009 | 3.524 |
| Gen-T | 0.976 | 0.867 | 0.004 | 1.326 | 0.976 | 0.867 | 0.004 | 1.326 | 0.971 | 0.807 | 0.004 | 1.490 |
Precision and Recall: Consider a Source Table and an output reclaimed table .
From the measure Tuple Difference Ratio (TDR) [24], we derive two similarity measures, Recall (Rec) and Precision (Pre), that measure the # of tuples in the intersection of and relative to the # of tuples in each table.
and .
In addition to metrics that measure the similarity between the tuples of a reclaimed table and a Source Table, we also include finer-grain metrics that measure the number of values that do not match within aligned tuples (tuples with the same key value). If there are multiple aligned tuples with respect to one tuple in the Source Table (multiple tuples in the reclaimed table with the same key value), then we consider the tuple that contains the largest number of column values shared with the corresponding tuple in the Source Table. This way, there is at most 1 aligned tuple in the reclaimed table for each tuple in the Source Table. In these measures, which we denote as divergence measures, the ideal score is 0 (the reclaimed table is identical to the Source Table). Specifically, we introduce Instance Divergence and Conditional KL-Divergence, enabling us to measure the nullified and erroneous values in the reclaimed table’s aligned tuples, with respect to the Source Table (see Section IV-A).
Instance Divergence: We measure how many missing values there are in each aligned tuple, with respect to its corresponding tuple in the Source Table. To do so, we use the inverse of Instance Similarity [20] (see Equation 2) and introduce Instance Divergence: .
Conditional KL-divergence: We aim to capture how many erroneous values there are in aligned tuples from a reclaimed table with respect to tuples in a Source Table. We adopt the traditional definition of conditional KL-divergence [82, 83], and add a penalization for erroneous values, such that the score is higher (diverges more) for reclaimed tables containing erroneous values as opposed to nulls in their aligned tuples with the Source Table (see the technical report [73] for more details). We report average scores over all Source Tables.
Efficiency Measures: For efficiency, we measure the average runtimes for all Source Tables, as well as the average ratio of the output reclaimed table’s size to the Source Table’s size (large integrations can significantly increase runtimes).
VI-B Effectiveness
| Method | Rec | Pre | Inst-Div. | D |
| ALITE | 0.704 | 0.128 | 0.095 | 1.332 |
| ALITE w/ int. set | 0.745 | 0.133 | 0.086 | 1.197 |
| ALITE-PS | 0.805 | 0.539 | 0.040 | 0.655 |
| ALITE-PS w/ int. set | 0.833 | 0.552 | 0.037 | 0.688 |
| Auto-Pipeline* | 0.674 | 0.272 | 0.158 | 2.574 |
| Auto-Pipeline* w/ int. set | 0.683 | 0.289 | 0.133 | 2.109 |
| Ver w/ int. set | 0.746 | 0.351 | 0.127 | 10.393 |
| Gen-T | 0.954 | 0.799 | 0.015 | 0.165 |
Tables III and II report the results for all methods on the TP-TR Small benchmark and the larger TP-TR benchmarks, respectively. For experiments on TP-TR Small, TP-TR Med, and TP-TR Large benchmarks, we input candidate tables discovered from just Set Similarity (Section V-A1). For experiments on SANTOS Large +TP-TR Med, we first discover relevant tables from the large data lake using Starmie [40], a state-of-the-art self-supervised system for scalable table discovery. Hence, it can discover a set of candidate tables for the Source Table from a large data lake. Although the primary use case of Starmie was table union search, it was shown to apply to other search semantics such as table discovery to improve the performance of downstream machine learning tasks via feature discovery (join search) and column clustering. Following Starmie, we run Set Similarity to find syntactically similar tables among the returned tables from Starmie.
We run baselines Auto-Pipeline* and Ver (w/ int. set) only on TP-TR Small benchmark (Table III), since they timeout for most, if not all Source Tables in the larger benchmarks. Ver times out when given the entire data lake from TP-TR Small. ALITE times out only on TP-TR Large benchmark (Table II). We discuss scalability and timeouts in Section VI-C.
Across all benchmarks, Gen-T outperforms the baselines for all metrics, while perfectly reclaiming 15-17 Source Tables across all benchmarks. Baselines ALITE-PS and Auto-Pipeline* only perfectly reclaim 3 Source Tables and 1 Source Table, respectively, across the benchmarks on which they do not time out, and ALITE and Ver does not perfectly reclaim any. In fairness, ALITE is an integration method that does not consider the Source Table (it is not ”target-driven” like Auto-Pipeline*). Also, Ver is a QBE method whose goal is to produce source tuples along with many additional tuples. In terms of similarity (Table II), Gen-T outperforms the top performing existing baseline method (ALITE-PS) by 11-25% in Recall and by 48-56% in Precision across all TP-TR benchmarks. For the divergence measures, we see that Gen-T produces tables that contain fewer nullified values in its aligned tuples with respect to the Source Table (Inst-Div.), as well as fewer erroneous values in its aligned tuples, which is reflected in the lower D scores than the baselines.
Even compared to each baseline that is given specified integrating sets of tables rather than large sets of candidates (‘w/ int. set’), Gen-T performs much better. Thus, the matrix traversal method (Section V-A3) used in Gen-T to refine the set of originating tables works well in filtering out misleading tables that could be integrated to produce tables containing erroneous values. We provide benchmark samples to exemplify this in our repository [80].
We now perform an analysis of the similarity measures for all methods on different types of queries used to form the Source Tables in TP-TR benchmarks, shown in Figure 6. Ranging from simple queries (that just perform Projection, Selection, and Union) to more complex queries (joining up to 3 tables and unioning up to 4 tables), we see that Gen-T outperforms the baselines on queries of all complexities used to initially create the Source Table. Thus, not only is the matrix traversal effective, but the set of operators used in table integration represents different types of queries well.
Tuning % Erroneous vs. Nullified Values: We further analyze Gen-T’s performance on data lake tables with different number of erroneous and nullified values in TP-TR Med tables (Figure 7). So far, TP-TR Med tables have 50% erroneous values in erroneous versions and 50% nulls in nullified versions (intersection point on the graph where Gen-T has 0.867 Precision). Now, we tune the percentage of values replaced with non-null, random strings (blue line in Figure 7) in erroneous versions, while the nullified versions always contain 50% nulls. Similarly, we tune the percentage of values replaced with nulls (red line in Figure 7) while holding the erroneous versions constant. For Gen-T to produce a perfect reclamation of a Source Table, it should only have originating tables with injected nulls so that these nulls can be replaced with correct values during table integration.
As data lake tables have more erroneous values (blue line), Gen-T is more likely to contain tables with nullified tuples in its set of originating tables, which results in an integrated table with higher precision. On the other hand, as we tune the percentage of values replaced with nulls (red line), precision decreases. As more nulls are injected, these tables also have fewer correct values. Gen-T is thus more inclined to have originating tables with 50% erroneous values, or 50% correct values, leading to a final integration with lower precision.
VI-C Scalability
Figure 8 shows the scalability of Gen-T, ALITE, ALITE-PS, and Auto-Pipeline* across benchmarks as the number and/or size of tables grows. Figure 8(a) reports average runtimes for all methods across all four benchmarks, starting from ingestion of the candidate tables. For Gen-T, this time includes the time it takes to prune the set of candidate tables to a set of originating tables, and integrate it to produce a reclaimed table. For other methods, this time only includes integration. We find that Auto-Pipeline* only runs on TP-TR Small without timing out, and ALITE, which performs full disjunction, is exponential in time and times out for the last two benchmarks. We set the timeouts as 30min for TP-TR Small, 7hrs for TP-TR Med and SANTOS Large +TP-TR Med and 24hrs for TP-TR Large.
Gen-T has a more consistent runtime across all benchmarks compared to all baselines. Gen-T is 3X faster compared to Auto-Pipeline* on TP-TR Small. On TP-TR Med, Gen-T is 40X faster than ALITE and on TP-TR Large, Gen-T is 5X faster than ALITE-PS. Thus, pruning candidate tables to originating tables seems to cut the cost of integration, a prevalent issue as shown by the baselines.
Figure 8(b) reports the average output sizes, or number of cell values in the reclaimed tables, with respect to the average Source Table sizes. As the number and size of tables grows across benchmarks, the output size relative to the size of the Source Table (expected output size) can easily grow at a fast rate if the integration is among more or larger tables, especially if it includes noisy tables from the real data lake (SANTOS Large). Output sizes for Gen-T remain consistent across all benchmarks (1.4-4.5X larger than the average Source Table size). This trend largely accounts for the higher precision of Gen-T since its output tables mostly consist of Source tuples. In contrast, output sizes for ALITE, ALITE-PS, and Auto-Pipeline* are 200-300X, 2.5-250X, and 4X larger than Source Tables’ sizes, respectively. Thus, Gen-T’s runtimes and output sizes remain consistent across benchmarks of different sizes.
VI-D Generalizability
We also experiment with the T2D Gold benchmark to see how well we can apply Gen-T to a real-world scenario with Web Tables. In this case, we do not know whether or how the tables were originally generated. Accordingly, we attempt to reclaim each table using a subset of other tables in the benchmark by iterating through each 515 tables as potential Source Tables. Gen-T successfully reclaims 3 Source Tables from an integration of multiple tables (5-6 tables), such that the outputs have perfect Recall, Precision, Instance-Div., and near-perfect D. Gen-T also finds duplicate tables for 12 Source Tables, or 6 sets of duplicates. This indicates that we can apply Gen-T in a different domain, even if no sources are known to be reclaimed, and retrieve successful reclamations. Baseline methods are able to reclaim 12-13 Source Tables, which are included in the 15 Source Tables reclaimed by Gen-T.
| Method | Recall | Precision | Inst-Div. | D |
| ALITE | 0.956 | 0.490 | 0.009 | 0.627 |
| ALITE-PS | 0.956 | 0.796 | 0.009 | 0.627 |
| Auto-Pipeline* | 0.881 | 0.725 | 0.088 | 19.261 |
| Gen-T | 0.956 | 1.000 | 0.009 | 0.627 |
We then run experiments on a data lake consisting of both T2D Gold and WDC Sample tables. This way, we evaluate how well the methods perform when a set of candidate tables returned from Set Similarity may contain irrelevant or misleading tables from the WDC Sample benchmark. Table IV presents the similarity and divergence scores for all methods on 33 of the common sources from T2D Gold for which all methods have non-empty, reasonably sized output tables. We can see that Gen-T outperforms the baselines for all measures, even having a precision of 1.0. In contrast, the baseline methods that are given the candidate tables from Set Similarity integrate all candidate tables and produce tables that contain many additional tuples.
VII Conclusion and Open Problems
Table Reclamation is essential in verifying if a data lake supports the tuples (facts) in a Source table. Our results show that, despite the large search space, Gen-T can solve the reclamation problem efficiently for source tables with keys. In future work, we will relax the key assumption with regard to source tables, and use a fast, approximate instance comparison algorithm to compare instances from a source table and data lake tables [84]. When a table can only be partially reclaimed, we plan to investigate whether the originating tables can be embedded in a new data lake and used to possibly generate a better reclamation. Alternatively, we plan to investigate if reclamation can be combined with data cleaning (for example, value imputation over missing values or entity resolution) to produce a better reclamation. In addition, we plan to consider the case in which values from a source table do not syntactically align with values from a data lake, in which case we can explore the semantic similarity of instances. Table reclamation can also be used to verify the tabular results of generative AI or large language models. Verifying the output of generative AI using a data management lens is an emerging and important area [85, 86]. For example, users who generate summary tables and charts (e.g. Microsoft Copilot [87]) or presentation slides (e.g. SlidesAI [88]) from input data would find it useful to verify model outputs and examine what data was used to generate them. Our approach will advance this area by allowing the automatic verification of tables that are created through complex integrations of other tables, something not yet considered in the literature.
Acknowledgements
This work was supported in part by NSF award numbers IIS-2107248, IIS-1956096, and IIS-2325632.
Appendix A Matrix Representations
A-A Expanding Candidate Tables
In order to represent candidate tables as matrices, their tuples need to align with those in the Source Table. However, not all candidate tables may share a key column with the Source Table. Thus, we need to join a given candidate table that does not share a key column with the Source Table with those that do. This way, tuples from all candidate tables can be aligned with tuples from a Source Table using key values.
As illustrated in Expand Algorithm 5, we traverse a graph that consists of candidate tables as nodes and we find a join path between candidate tables that do not have a Source Table’s key (start nodes), and candidate tables that do (potential end nodes). If two candidate tables can join on common columns, their nodes are connected with an edge. For each edge, we use standard join-size cardinality estimation to find edge weights [71].
After we find a path from a start node to an end node, we iteratively join all tables in the path, resulting in a table that shares a key column with the Source Table. This way, all candidate tables share a key column with the Source Table.
A-B Two-Valued vs. Three-Valued Matrix Representations
After aligning tuples in candidate tables to Source Table’s tuples that share the same key values, we can represent candidate tables as matrices to show how similar the values in candidate tables are to those in the Source Table. These matrices have the same dimensions and indices as the Source Table. First, we consider matrices that consist of binary values, where a 0 in tuple , column represents a value at index () in the candidate table that is different from the value in the same position in the Source Table, and a 1 represent common values in the same indices.
However, populating alignment matrices with binary values does not fully encode how many values in candidate tables differ from those in the Source Table. Specifically, this representation does not distinguish between nullified values in candidate tables (null values in candidate tables at index () where there is a non-null value in the Source Table at the same position), and erroneous values (different non-null values in candidate tables at index () from those in Source Tables at the same position).
Instead, we encode matrix representations using three values, where at a given index in an aligned tuple, there is a 1 for a value shared between a candidate table and the Source Table, 0 for a null value in the candidate table where there is a non-null value in the Source Table, and -1 for a non-null value in the candidate table that differs from the value in the Source Table.
A-C Metrics
Conditional KL-divergence: Given column shared between a Source Table and a reclaimed table , suppose we have probability distributions, for in the Source Table and for in the reclaimed table. We condition on the key values in key column . The conditional KL-divergence (or conditional relative entropy) between and of sample space of column conditioned on key is as follows:
Given non-key columns in a Source Table we take the average D for each column divided by the probability of a key value in matching a key value from the Source Table () and the number of non-key columns (). Then, the conditional KL-divergence of the reclaimed table is as follows:
The conditional KL-divergence of the reclaimed table is a score , with 0 being the ideal score. There is no upper limit on this metric since it naturally approaches when no key value from the Source Table is found in the reclaimed table.
A-D Effectiveness Experiments
Effectiveness of LLM Baseline: To further assess the effectiveness of Gen-T, we compare it to OpenAI’s ChatGPT3.5 [89] as a representative Large Language Model (LLM). In our prompt for ChatGPT, we define the table reclamation problem and input a source table and a set of tables. We ask ChatGPT to return an integration result that reproduces the source table as closely as possible. Due to scalability constraints, we only use the integrating set as input.
In our effectiveness experiments, we run ChatGPT on TP-TR Small benchmark, since it times out for all source tables in the larger benchmarks. We find that ChatGPT achieves a Recall of 0.239 and a precision of 0.256. For divergence measures, ChatGPT achieves an Instance-Divergence of 0.540 and a D of 209.83. ChatGPT returns tables that only contain some source tuples, while containing many non-null erroneous values (high D score).
Effectiveness of Pruning in Gen-T: For a more detailed analysis, we analyze Recall, Precision, and F1-Scores of Gen-T and baseline, ALITE-PS, on TP-TR Med for each of its 26 Source Tables (Figure 9). Note that ALITE-PS directly integrates a set of candidate tables, whereas Gen-T first prunes the set of candidate tables to a set of originating tables before performing table integration. Gen-T outperforms ALITE-PS in Precision for all Source Tables, and outperforms ALITE-PS in Recall for 24 of 26 total Source Tables. This shows that ALITE-PS, which directly integrates candidate tables without pruning, reclaims more Source Tuples than Gen-T, which does include a pruning step, for only a few Source Tables. Also, Gen-T outperforms ALITE-PS in F1-Score for all Source Tables (Figure 9(c)), showing that even if ALITE-PS outperforms Gen-T in Recall, it does not impact the F1-Score.